HUBBLE SPACE TELESCOPE DAILY REPORT #2380 PERIOD COVERED: 0000Z (UTC) 05/25/99 - 0000Z (UTC) 05/26/99 Daily Status Report as of 146/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled with anomalies. 1.2 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled with no problems. 1.3 Completed Two Sets of STIS/CCD 7949 (CCD Dark and Bias Monitor -- Cont. to June 99) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a CCD dark calibration and a bias monitoring program. It was also scheduled to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Four Sets of WF/PC-2 7282 (A Continuing Snapshot Survey for Compact Stellar Groups and Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to make observations of the Wolf-Rayet stars WR138, CYGOB2N09, WR92 and WR98 as part of a continuing snapshot survey of compact stellar groups and clusters. The newly imaged clusters allow us to accurately determine the distances, luminosities, and ages of the WR members. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 8290 (Weather, Moons, and Orbit of the Brown Dwarf Gl 229B) The WF/PC-2 was used to make observations of the brown dwarf star HD42581 (a.k.a., GI 229B). Gl 229B is the only unambiguous brown dwarf. It can be imaged with high astrometric and photometric accuracy only by HST. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Five Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled with no anomalies. 1.7 Completed WF/PC-2 7470 (A Snapshot Survey of Galactic Globular Clusters) The WF/PC-2 was used to make snapshot survey observations of the crowded center of the galactic globular cluster NGC-6355. The resulting color-magnitude diagram will include all stars down to about one magnitude below the main-sequence turnoff. The observations were completed as scheduled with no problems. 1.8 Completed WF/PC-2/STIS/CCD 7615 (Determination of the Mass Function of a Globular Cluster Through Gravitational Microlensing) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to make gravitational microlensing observations of the large globular cluster NGC-6656 located in Sagittarius in order to study its mass function. Gravitational microlensing offers a unique and direct method of determining the masses of stars and other compact objects down to planetary masses. A bias and dark calibration observations were also taken. The observations were completed as planned, and no problems were reported. 1.9 Completed WF/PC-2 7378 (Narrow Band Imaging of Jets and Bipolar Outflows From Symbiotic Stars) The WF/PC-2 was used to make observations of the symbiotic star HD182917 to make narrow band imaging of its jets and bipolar outflows. With these HST observations we have the unique opportunity to follow in real-time the spatial evolution of the outflows {on a time basis of very few years}, and therefore to derive 3-D dynamical models of their structure. The observations were completed as scheduled, and no problems were reported. 1.10 Completed STIS/CCD 8470 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition VI proposal. The observations were completed as scheduled with no anomalies. 1.11 Completed WF/PC-2 8210 (HST Imaging of Bi-Polar Proto-Planetary Nebulae and Circumstellar Arcs) The WF/PC-2 was used to make observations of the bi-polar proto-planetary nebula IRAS20136. Proto-planetary nebulae provide the best opportunity to test theories of asymptotic giant branch mass loss {symmetric/asymmetric, continuous/episodic} and the processes which shape planetary nebulae. The observations were completed as planned, and no problems were reported. 1.12 Completed STIS/CCD/MA1 7437 (Understanding High-Redshift and Starburst Galaxies: A UV Spectroscopic Survey of O-Stars in the SMC) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make ultraviolet spectrographic observations of galaxy AV83. The data from this survey, of importance in many areas of astrophysics, will immediately be made public as no proprietary period is requested. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 7 Successful : 7 As documented in HSTAR 7147, there was an FGS loss of lock at 146/044449Z. Two proposals occurring during the period of the next report were affected: WF/PC-2 8059 and STIS 7437. 2.2 FHST Updates: Scheduled : 22 Successful : 22 2.3 Operations Notes: The Flight Operations Team is in the SAMS back-up facility, using the PRS ground system. GSFC Internal Simulation #4 (EVA-3/Deploy) will be conducted today and tomorrow, running from 8 a.m. until approximately 4:00 p.m. each day. The SSR EDAC error counter was cleared four times, using ROP SR-1. The NICMOS FPA temperature sensor for detector #3 was enabled at 1446/0549Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH